OpenStudy (anonymous):

Betty and Karen have been hired to paint the houses in a new development. Working together the women can paint a house in two-thirds the time that it takes Karen working alone. Betty takes 6h to paint a house alone. How long does it take Karen to paint a house working alone?

hero (hero):

In general, to find the time it takes two people, B and K to work together, use the following formula: \(\large \frac{BK}{B + K} = t\) where B = Betty K = Karen t = time it takes when Betty and Karen are working together

hero (hero):

Adding B = 6 and \(t = \frac{2K}{3}\) to the formula we get: \[\large \frac{6K}{6 + K} = \frac{2K}{3}\]

hero (hero):

Now, just cross multiply and continue solving for K

hero (hero):

3(6K) = 2K(6 + K) 18K = 12K + 2K^2 0 = 12K - 18K + 2K^2 0 = -6K + 2K^2 0 = 2K^2 - 6K 0 = 2K(K - 3) 2K = 0 K - 3 = 0 K = 3 Therefore, it takes Karen 3 hours to paint a house while working alone.